Minutes of the 4th Committee Meeting of the Dublin University Rifle Club

Moderator Steven Watterson

Minutes taker Mark Dennehy

Present Steven Watterson (Captain), Mark Dennehy (Secretary/Match Secretary), John Keeney (Ammo&Equipment Officer),

Absent(with apologies) Benjamin Jones (Treasurer), Alan Smyth (Chief Range Officer),

Absent Conor McDermottroe (Public Relations Officer), Jason Sibley, Deirdre Ninaber (Ladies’ Coordinator)

Date 2 July 2001 1800 – 1930

Distribution DURC Committee

Overview of topics

1 Minutes 1

2 Junior Squad 1

3 Ammo&Equipment officer’s report 2

4 Treasurers Report 2

5 Secretarys Report 2

6 Matters arising from Minutes 3 6.1 Meeting Times . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 6.2 Calendar . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 6.3 Medals replacing perpetual trophies . . . . . . . . . . . . . . . . . . . 3

1 Minutes

6.4 3P Air . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3

7 DURC 50m Open 3

8 Conor and Peter’s Training 5

9 Next Meeting 5

1 Minutes

The minutes were accepted from the previous meeting, but Mark has still to reformat them in the normal manner - this to be presented at the next meeting. Minutes from the 2nd meeting were formally signed.

2 Junior Squad

Liam Murray cannot try for the Junior Squad, being in Galway for the remainder of the summer. Conor and Peter are the only two left who can try out and they need more training.

3 Ammo&Equipment officer’s report

John reported that RD Sports (the main Eley supplier in Ireland) will sell us Eley ammunition in batches of 5000, at the following prices :

Tenex £140 per 1000 plus VAT
Match £84 per 1000 plus VAT

Mark asked about EPS eley ammunition, John reported that RD have it but can’t seem to sell it - they have batches but have only sold four to date. Mark asked about the size but 5000 rounds was too much. Steven pointed out that we really don’t need EPS, Mark agreed but said he’d still like to try it.

John reported that the planned delivery of Lapua Master, T22 and Air pellets had not taken place on Friday June 30th as planned but would happen before the Nationals on the 7th. Task: Get ammunition delivered to college before the Nationals (John)

John has still to talk to Pat Moray and Tim Cooper on Wednesday morning to sort out the ammunition accounting procedures and storage arrangements regarding the new security offices. Steven suggested a locker in the new offices was the best approach. John agreed, but pointed out that Pat Moray’s concern was the size of his safe. Mark said that the best idea was a bigger safe and John agreed as he would like to see us buying ammunition in batches of 10,000 and buying several batches of different ammunition types. Task: Meet with Pat Moray and Tim Cooper about ammunition arrangements (John)

4 Treasurers Report

John reported that Rifle #4 was giving problems again, and that when he’d gone to clean the bolt it was full of gunk. Mark pointed out that this had happened earlier this year and that while he’d been able to clean it and get it to work, he had stated then that it was a temporary fix and that we needed to get Dave Cooney to service it.

John reported ammunition levels of a few hundred rounds per type, with more due in mid-week. All Super Club had to be shot by wednesday to facilitate meeting with Pat Moray & Tim Cooper. Steven asked how the Super Club matter would be handled by John, John reported that he had good working relations with Pat Moray and that he did not anticipate a crisis.

4 Treasurers Report

There was no report, the Treasurer being absent, but there was no reason to expect an improvement in our fiscal situation.

5 Secretarys Report

Mark reported the sending out of CARs for the FSC Open, and that CARs were being prepared for the training weekend in FSC and the nationals. John worried that these reports were reaching spam mail levels. Mark replied that this was fine, so long as the people involved had the name of DURC in their heads as attending competitions and getting results.

Mark also reported receiving a catalog and pricelist from Edelmann targets. John wondered if we needed to get these from someone other than Dave Cooney, Mark replied that he had gotten them to provide alternative avenues for purchasing our expendables in case Dave took the loss of business from our ammo to be sufficent reason not to deal with us. John though this unlikely, Mark pointed out that price hikes were not unlikely.

6 Matters arising from Minutes

6.1 Meeting Times

Mark was happy enough with Monday meetings but insisted that they be held in the evening as lunchtime was not a good time due to pressures from classes and so forth. John said that he didn’t like 1800 as he was just starting Counter-Strike games at that point. Steven suggested 1900 - Mark said this gave people a chance to finish lectures, get some food and prepare for the meeting, John thought undergrads might not like the time. Steven suggested we try it and see. This was agreed on.

6.2 Calendar

Mark presented the latest draft of the Irish Marksman with the current NTSA calendar and squad times, and suggested that we could now set the DURC calendar, but should wait for a quorum to do so.

7 DURC 50m Open

6.3 Medals replacing perpetual trophies

Mark agreed that we had no money to buy a medal this year, but that we had to find money for this next year as it was too cheap to do otherwise.

6.4 3P Air

Mark asked if we were using a rubber catchment sheet for the catchers - Steven pointed out that we didn’t use one for the standing catchers, so why bother?

7 DURC 50m Open

The scheduling of the DURC 50m Open was considered. After looking at various weekends and conflicts, including committee members birthdays, the weekend of July 21-22 was decided on. Logistics were then discussed and narrowed down to the following areas: Targets. It was agreed that Ben should chase down the targets we had allready ordered from the NSRA and get them delivered post-haste. The 100yd targets were to be bought from Dave Cooney if we went ahead with the Scottish Match. Task: Chase down 50m Targets bought from the NSRA (Ben) Task: Cost 100 100yd 3-bull targets from Dave Cooney (John)

People. Mark is to advertise the competition in the Irish Marksman and send out notice letters. Task: Advertise DURC Open in Irish Marksman (Mark) Task: Send out DURC Open notice letters to clubs (Mark)

Barbecue. Mark has notes on last years barbecue and the lessons learnt from it and will handle the logistics for this. There were queries regarding vegetarian cooking at a barbecue, Steven said that it was difficult but grilled vegetables were about the best that could be done. Task: Learn how to barbecue vegetables (Mark) Task: Arrange logistics of Barbecue for Open (Mark)

Scoring. It was agreed to ask Keith if he would score and Mark would supply the laptop computer as soon as he had it fixed. Task: Ask Keith to score for DURC Open (Unassigned) Task: Fix laptop computer (Mark)

Range Hire. John agreed to talk to Dave Cooney about getting the FSC range for that weekend, as well as the 100yd targets. Task: Arrange FSC range for the DURC Open with Dave Cooney (John)

Squaddings and Timing. A long discussion on timings took place to allow a 100yd and 50m shoot on the same day. The early start was lamented by John, but Mark and Steven insisted that it was normal. Mark suggested that he might be able to store gear in his house to save travel time as he was only 4 miles from FSC and not 40 as DURCs range was. This had to be cleared by his parents however. The final arrangements were to be half-hour details for squadding, starting at 0930 and running until 1400, with the

8 Conor and Peter’s Training

Scottish Match taking place at 1500, again with 30 minute details. This 30-minute detail included target changing. Prize-giving would follow the Scottish Match.

Prizes. 18 prizes will be needed, 15 for the English Match – Championship (3), Classes A,B,C,D (3 each) – and 3 for the Scottish Match. It was suggested that a Junior section be included, but this was ruled out. Glassware and “prize-type prizes” were suggested for classes D and C where it was important to the people to get medals and so forth. This was agreed. Championship prizes also had to be offical-type prizes, but classses A and B were agreed to be of a novel nature, though noone had any ideas as to what that should be. Mark suggested we consider it and discuss via email. Task: Decide on prizes for Open (All)

Fees. Fees were discussed but not agreed on, though the ballpark figures seemed to be £12 for the English Match and £8 for the Scottish Match - what the Barbecue costs added to this was unknown and this must be discussed more later.

Steven wondered if there would be a final, but it was decided not to have one due to the difficulty of scoring one. The prospect of a Scottish Match was discussed and the general feeling was that it was a good idea and we should go ahead with it.

8 Conor and Peter’s Training

John was worried that Conor and Peter were not training enough. Steven disagreed, he thought they were putting in the hours but just didn’t know how to train, and that the onus was on us to provide that. Mark suggested that we simply had no formal training techniques - no exercises, no goal scores set, etc. This was agreed and it was agreed also that we would supervise them more closely with this in mind from now on. John also agreed to ask Dave Cooney if they could shoot for classification in the Nationals on the Friday as they couldn’t compete on the weekends due to work committments. Task: Ask Dave Cooney about early shooting in Nationals for classification (John)

9 Next Meeting

The next meeting was set for 1900 on Monday July 9th, in the ORI, either in the Common Room or the origf printer room if that proved too hot for John. Steven reported he would be absent from the 5th to the 19th for the Island Games, Mark reminded him to shoot his cards for the first round of the Trinity Top Ten while there, and that he would get the stickers to him before he left.

T.B.N.O.B.T.M.W.D.C.

Steven Watterson, Club Captain 9 Next Meeting

Mark Dennehy, Club Secretary

9 Next Meeting

Tasks agreed on at meeting

John Get ammunition delivered to college before the Nationals (2 July

2001) [/157] . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . /2

John Meet with Pat Moray and Tim Cooper about ammunition arrange-

ments (2 July 2001) [/170] . . . . . . . . . . . . . . . . . . . . . . . . . /2

Ben Chase down 50m Targets bought from the NSRA (2 July 2001) [/250] /2 John Cost 100 100yd 3-bull targets from Dave Cooney (2 July 2001) [/251] /2 Mark Advertise DURC Open in Irish Marksman (2 July 2001) [/258] … /2 Mark Send out DURC Open notice letters to clubs (2 July 2001) [/259] . . /2 Mark Learn how to barbecue vegetables (2 July 2001) [/268] . . . . . . . . /2 Mark Arrange logistics of Barbecue for Open (2 July 2001) [/269] . . . . . /2 Unassigned Ask Keith to score for DURC Open (2 July 2001) [/275] . . . . . . . /2 Mark Fix laptop computer (2 July 2001) [/276] . . . . . . . . . . . . . . . /2 John Arrange FSC range for the DURC Open with Dave Cooney (2 July

2001) [/283] . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . /2

All Decide on prizes for Open (2 July 2001) [/310] . . . . . . . . . . . . /2 John Ask Dave Cooney about early shooting in Nationals for classification

(2 July 2001) [/340] . . . . . . . . . . . . . . . . . . . . . . . . . . . . . /2